OpenStudy (smurfy14):

((sqrt of -2) + 2i)(6+5i)

OpenStudy (smurfy14):

\[(\sqrt{-2}+2i)(6+5i)\]

OpenStudy (anonymous):

first turn it in to \[(\sqrt{2}i+2i)(6+5i)\] then do 4 multiplications

OpenStudy (anonymous):

\[6\sqrt{2}i+12i+5\sqrt{2}i^2+10i^2\] then say "oh, i^2=-1\] and get \[6\sqrt{2}i+12i-5\sqrt{2}-10\]

OpenStudy (anonymous):

\[-5\sqrt{2}-10+(6\sqrt{2}+12)i\] in standard form

OpenStudy (smurfy14):

oh ok thanks so much satellite!

OpenStudy (anonymous):

yw